# From graph to Riesz continuity

### v2

@@ 5 Self-adjoint operators
$p_{\lambda}=\operatorname{\mathbb{1}}_{[\lambda,+\infty)}\colon\mathcal{R}^{\mathrm{sa}}_{\lambda}\to\mathcal{P}(H),\quad\mathcal{R}^{\mathrm{sa}}_{\lambda}=\left\{A\in\mathcal{R}^{\mathrm{sa}}(H)\,|\,\lambda\in\operatorname{Res}(A)\right\}$
[…]
@@
Suppose that $\mathcal{A}$ is Riesz continuous. Then the family $a_{x}=\bm{\chi}(\mathcal{A}_{x})$ is norm continuous and $\mu=\bm{\chi}(\lambda)$ lies in the resolvent set of each $a_{x}$. Therefore, the map $x\mapsto p_{\lambda}(x)=\operatorname{\mathbb{1}}_{[\mu,+\infty)}(a_{x})$ is also norm continuous.
Conversely, suppose that $p_{\lambda}$ is norm continuous. Then $r_{\lambda}=2p_{\lambda}-1$ is a norm continuous family of symmetries commuting with $\mathcal{A}$. The family $\mathcal{A}^{\prime}=\mathcal{A}\cdot r_{\lambda}$ of self-adjoint operators is graph (resp Riesz) continuous if and only if $\mathcal{A}$ is graph (resp Riesz) continuous. But, in contrast with $\mathcal{A}$, the family $\mathcal{A}^{\prime}$ is uniformly bounded from below by $-|\lambda|$. The graph and Riesz topology coincide on the set of operators bounded from below by a given constant [$P_{2}$, Proposition 11.3], so $\mathcal{A}^{\prime}$ and thus also $\mathcal{A}$ are Riesz continuous. $\square$
Polarization associated with operator family.   Let $\mathcal{A}$ be a graph continuous family of self-adjoint operators with compact resolvents acting on fibers of a Hilbert bundle $\mathcal{H}$. Then $X_{\lambda}$ is open for every $\lambda\in\mathbb{R}$ by [Ka, Theorem IV.3.1] and these sets form an open covering of $X$. By Proposition 5.2, $p_{\lambda}$ is strongly continuous in local trivializations of $\mathcal{H}$, so it determines a decomposition of $\mathcal{H}$ over $X_{\lambda}$ into the orthogonal sum $\mathcal{H}_{\lambda}^{+}\oplus\mathcal{H}_{\lambda}^{-}$ of two Hilbert subbundles. For $\lambda<\mu$ the difference (5.2) is a norm continuous family of finite rank operators over $X_{\lambda}\cap X_{\mu}$ by [Ka, Theorem IV.3.16].
Suppose additionally that each $\mathcal{A}_{x}$ is neither essentially positive nor essentially negative. Then $p_{\lambda}(x)\in\mathcal{P}^{\star}(\mathcal{H}_{x})$, so the collection $(X_{\lambda},p_{\lambda})$ is an atlas of a polarization. Hence a family $\mathcal{A}$ satisfying assumptions above determines a polarization of $\mathcal{H}$ which we call the polarization associated with $\mathcal{A}$ and denote by $\Pi_{\mathcal{A}}$. (The associated polarization $\Pi_{\mathcal{A}}$ is actually finite, but we are interested only in its equivalence class with respect to compact deformations, so we will consider it as a compact polarization.)
@@
More generally, consider a family $\mathcal{A}$ of Fredholm self-adjoint operators acting on fibers of $\mathcal{H}$. Suppose that $\mathcal{A}$ is graph quasi-continuous (that is, $\mathcal{A}$ admits local trivializations taking it to a graph continuous family). Then the associated polarization $\Pi_{\mathcal{A}}$ is defined in a similar manner. The only change is in the definition of the set $X_{\lambda}$: one should add the condition that the interval $[\lambda,0]$ or $[0,\lambda]$, depending on the sign of $\lambda$, is disjoint from the essential spectrum of $\mathcal{A}_{x}$.
If a trivialization is adapted to such a family $\mathcal{A}$, then it is adapted to $\Pi_{\mathcal{A}}$ as well. However, the second direction of Proposition 5.4 is no longer true: a trivialization adapted to $\Pi_{\mathcal{A}}$ is not necessarily adapted to $\mathcal{A}$, because it might turn $\mathcal{A}$ into a graph discontinuous family and then Proposition 5.3 cannot be applied. This is the place where condition of compact resolvents comes into the picture: the graph continuity of a family of operators with compact resolvents is independent of the choice of a trivialization.
Proof of Theorem E.   Let $\Pi_{\mathcal{A}}$ be the polarization of $\mathcal{H}$ associated with $\mathcal{A}$. By Proposition 5.4, a (local or global) trivialization of $\mathcal{H}$ is adapted to $\mathcal{A}$ if and only if it is adapted to $\Pi_{\mathcal{A}}$. It remains to apply Theorem 4.4. $\square$
Example 5.5.
Let us show that Theorem E does not hold without additional assumptions on the base space $X$. We will construct a graph continuous family $\mathcal{A}\colon X\to\mathcal{R}^{\star}_{K}(H)$ of invertible self-adjoint operators with compact resolvents over a compact metrizable space $X$ such that $\mathcal{A}$ is not Riesz quasi-continuous. Moreover, for every point $x\in X$, $\mathcal{A}$ is not Riesz quasi-continuous at $x$ (that is, it is not Riesz quasi-continuous in any neighborhood of $x$).
Our example is based on a construction of a non-trivial separable Hilbert bundle $\mathcal{E}$ over a compact space $X$ by Dixmier and Douady. See [DD, Theorem 6]. Let us recall their construction. Let $H_{i}=\mathbb{C}^{2}$, $X_{i}=\mathbb{P}(H_{i})\cong\mathbb{C}\mathbb{P}^{1}$, $H=\oplus_{i=1}^{\infty}H_{i}$, and $X=\prod_{i=1}^{\infty}X_{i}$ with the product topology. The space $X$ is a countable product of metrizable spaces and thus is metrizable itself; it is a product of compact spaces and thus is compact itself. The Hilbert bundle $\mathcal{E}$ over $X$ is defined as the subbundle of the trivial Hilbert bundle $H_{X}$ with the fiber $\mathcal{E}_{x}=\oplus_{i=1}^{\infty}x_{i}$ over $x=(x_{i})$, where $x_{i}\in X_{i}$ is considered as a one-dimensional subspace of $H_{i}$. By [DD, Theorem 6], every section of $\mathcal{E}$ vanishes at some point. In particular, $\mathcal{E}$ is not trivial and thus it is not locally trivial (since all locally trivial Hilbert bundles over paracompact base spaces are trivial).
Moreover, the bundle $\mathcal{E}$ looks the same near each point of $X$ (more precisely, the group $\prod_{i=1}^{\infty}\mathcal{U}(H_{i})$ acts transitively on $X$ and this action lifts to the action on $\mathcal{E}$ by bundle automorphisms), so a local triviality of $\mathcal{E}$ near one point would imply local triviality over the whole $X$. Therefore, for every open subset $X^{\prime}\subset X$, $\mathcal{E}$ is not trivial over $X^{\prime}$.
Let $c\colon\mathbb{N}\to\mathbb{R}_{>0}$ be a sequence of positive numbers converging to $\infty$. Then
$A=\oplus_{i=1}^{\infty}c_{i}\cdot\operatorname{Id}_{H_{i}}$
is a positive regular operator with compact resolvents. Let $p(x)=\oplus_{i=1}^{\infty}p_{i}(x)$ be the orthogonal projection of $H$ onto $\mathcal{E}_{x}$, and let $r(x)=2p(x)-1$ be the corresponding symmetry. Then $r(x)$ commutes with $A$, and $\mathcal{A}(x)=A\cdot r(x)$ is an invertible self-adjoint operator with compact resolvents, which is neither essentially positive nor essentially negative.
We claim that the family $\mathcal{A}\colon X\to\mathcal{R}^{\star}_{K}(H)$ is graph continuous. Indeed, $(1+\mathcal{A}_{x}^{2})^{-1}=(1+A^{2})^{-1}$ is a compact operator independent of $x$. The map
$x\longmapsto\mathcal{A}_{x}(1+\mathcal{A}_{x}^{2})^{-1}=r_{x}A(1+A^{2})^{-1}$
is the product of a strongly continuous family of symmetries $r_{x}$ and a fixed compact operator $A(1+A^{2})^{-1}=\bm{\chi}(A)(1+A^{2})^{-1/2}$ and thus is norm continuous. Therefore, all components of the orthogonal projection (2.6) onto the graph of $\mathcal{A}_{x}$ are norm continuous and thus $\mathcal{A}$ is graph continuous.
However, for every open subset $X^{\prime}\subset X$, the restriction of $\mathcal{A}$ to $X^{\prime}$ is not Riesz quasi-continuous. Indeed, the positive spectral projection $p(x)$ of $\mathcal{A}(x)$ is the orthogonal projection of $H$ onto $\mathcal{E}_{x}$. If $\mathcal{A}$ were Riesz continuous in some local trivialization of the trivial Hilbert bundle $X^{\prime}\times H$, then the map $x\mapsto p(x)$ would be norm continuous in the same local trivialization and $\mathcal{E}$ would be a locally trivial Hilbert bundle over $X^{\prime}$, which is not the case.
Remark.
@@
If one multiplies the family $\mathcal{A}_{x}$ from this example by the strongly continuous family of unitary operators $r_{x}$, then the resulting family $x\mapsto\mathcal{A}_{x}\cdot r_{x}=A$ will be Riesz continuous (and even constant). However, for every strongly continuous family of unitary operators $v_{x}\in\mathcal{U}(H)$, the family $x\mapsto v_{x}\mathcal{A}_{x}v_{x}^{*}$ is Riesz discontinuous at every point $x\in X$. This demonstrates the difference between Theorems C and D, that is, between one-sided multiplication and conjugation by unitary operators, which we discussed in the Introduction.

### v3

@@ 7 Self-adjoint operators
$p_{\lambda}=\operatorname{\mathds{1}}_{[\lambda,+\infty)}\colon\mathcal{R}^{\mathrm{sa}}_{\lambda}\to\mathcal{P}(H),\quad\mathcal{R}^{\mathrm{sa}}_{\lambda}=\left\{A\in\mathcal{R}^{\mathrm{sa}}(H)\mid\lambda\in\operatorname{Res}(A)\right\}$
is graph-to-strong continuous. Let $A_{0}\in\mathcal{R}^{\mathrm{sa}}_{\lambda}$. Then $\operatorname{Res}(A_{0})$ contains the interval $\lambda-\varepsilon,\lambda+\varepsilon]$ for some $\varepsilon>0$. By [Kat, Theorem IV.3.1], there is a neighborhood $V$ of $A_{0}$ in ${{}^{g}\mathcal{R}}^{\mathrm{sa}}$ consisting of operators whose resolvent set contains this interval. Let $f\colon\mathbb{R}\to[0,1]$ be a continuous function equal to 0 on $(-\infty,\lambda-\varepsilon]$ and to 1 on $\lambda+\varepsilon,\infty)$. The function $g=\bm{\chi}-(2f-1)$ is continuous on $\mathbb{R}$ and vanishes at $\pm\infty$, so it can be approximated by polynomials in $(x+i)^{-1}$ and $(x-i)^{-1}$ and thus induces a graph-to-norm continuous map $g\colon\mathcal{R}^{\mathrm{sa}}\to\mathcal{B}$. See [RS], Theorem VIII.20(a). By Proposition 2.1 $\bm{\chi}$ is graph-to-strong continuous, so $f=(\bm{\chi}-g+1)/2$ is also graph-to-strong continuous. Since $p_{\lambda}$ coincides with $f$ on $V$ and $A_{0}$ was chosen arbitrary, $p_{\lambda}$ is graph-to-strong continuous on the whole $\mathcal{R}^{\mathrm{sa}}_{\lambda}$. $\square$
@@
Suppose that $\mathcal{A}$ is Riesz continuous. Then the family $a_{x}=\bm{\chi}(\mathcal{A}_{x})$ is norm continuous and $\mu=\bm{\chi}(\lambda)$ lies in the resolvent set of each $a_{x}$. Therefore, the map $x\mapsto p_{\lambda}(x)=\operatorname{\mathds{1}}_{[\mu,+\infty)}(a_{x})$ is also norm continuous.
Conversely, suppose that $p_{\lambda}$ is norm continuous. Then $r_{\lambda}=2p_{\lambda}-1$ is a norm continuous family of symmetries (that is, self-adjoint unitary operators) commuting with $\mathcal{A}$. By Proposition 2.2, the family $\mathcal{A}^{\prime}=\mathcal{A}\cdot r_{\lambda}$ of self-adjoint operators is graph (resp Riesz) continuous if and only if $\mathcal{A}$ is graph (resp Riesz) continuous. But, in contrast with $\mathcal{A}$, the family $\mathcal{A}^{\prime}$ is uniformly bounded from below by $|\lambda|$. The graph and Riesz topology coincide on the set of operators bounded from below by a given constant [P2, Proposition 11.2], so $\mathcal{A}^{\prime}$ and thus also $\mathcal{A}$ are Riesz continuous. $\square$
Polarization associated with operator family.  Let $\mathcal{A}$ be a graph continuous family of self-adjoint operators $\mathcal{A}_{x}\colon\mathcal{H}_{x}\to\mathcal{H}_{x}$ with compact resolvents. Then $X_{\lambda}$ is open for every $\lambda\in\mathbb{R}$ by [Kat, Theorem IV.3.1] and these sets form an open covering of $X$. By Proposition 7.1, $p_{\lambda}$ is a bundle projection of $\mathcal{H}$ over $X_{\lambda}$. For $\lambda<\mu$ the difference (7.2) is a norm continuous family of finite rank operators over $X_{\lambda}\cap X_{\mu}$ by [Kat, Theorem IV.3.16]. Suppose additionally that each $\mathcal{A}_{x}$ is neither essentially positive nor essentially negative. Then the collection $(X_{\lambda},p_{\lambda})$ is an atlas of a (finite) polarization. Hence a family $\mathcal{A}$ satisfying assumptions above determines a polarization of $\mathcal{H}$ which we call the polarization associated with $\mathcal{A}$ and denote by $\Pi_{\mathcal{A}}$.
@@
Proof of Theorem 7.  Let $\Pi_{\mathcal{A}}$ be the polarization of $\mathcal{H}$ associated with $\mathcal{A}$. By Proposition 7.3, a (local or global) trivialization of $\mathcal{H}$ is adapted to $\mathcal{A}$ if and only if it is adapted to $\Pi_{\mathcal{A}}$. It remains to apply Theorem 5. $\square$
Remark 7.4.
More generally, one can consider a family $\mathcal{A}$ of Fredholm self-adjoint operators acting on fibers of $\mathcal{H}$. Suppose that $\mathcal{A}$ admits local trivializations taking it to a graph continuous family. Then the associated polarization $\Pi_{\mathcal{A}}$ is defined in a similar manner. The only change is in the definition of the subset $X_{\lambda}$: one should add the condition that the interval $\lambda,0]$ or $0,\lambda]$, depending on the sign of $\lambda$, is disjoint from the essential spectrum of $\mathcal{A}_{x}$.
If a trivialization is adapted to such a family $\mathcal{A}$, then it is adapted to $\Pi_{\mathcal{A}}$ as well. However, the converse statement is no longer true: a trivialization adapted to $\Pi_{\mathcal{A}}$ is not necessarily adapted to $\mathcal{A}$, because it might turn $\mathcal{A}$ into a graph discontinuous family and then Proposition 7.2 cannot be applied. This problem arises already for continuous maps $X\to{{}^{g}\mathcal{R}}^{\star}_{F}(H)$. This is the place where condition of having compact resolvents comes into the picture: the graph continuity of a family of operators with compact resolvents is independent of the choice of a trivialization.
Example 7.5.
Let us show that Theorem 7 does not hold without additional assumptions on the base space $X$. We will construct a graph continuous family $\mathcal{A}\colon X\to\mathcal{R}^{\star}_{K}(H)$ of invertible self-adjoint operators with compact resolvents over a compact metrizable space $X$ such that, for every point $x\in X$, $\mathcal{A}$ has no adapted trivialization in any neighborhood of $x$.
Our example is based on a Dixmier–Douady’s construction of a non-trivial separable continuous field $\mathcal{E}$ of Hilbert spaces over a compact space $X$. See [DD, Theorem 6]. Let us recall their construction. Let $H_{i}=\mathbb{C}^{2}$, $X_{i}=\mathbb{P}(H_{i})\cong\mathbb{C}\mathbb{P}^{1}$, $H=\oplus_{i=1}^{\infty}H_{i}$, and $X=\prod_{i=1}^{\infty}X_{i}$ with the product topology. The space $X$ is a countable product of metrizable spaces and thus is metrizable itself; it is a product of compact spaces and thus is compact itself.
For $x=(x_{i})\in X$, let $\mathcal{E}_{x}\subset H$ be the closed span of $\oplus_{i=1}^{\infty}x_{i}$, where $x_{i}\in X_{i}$ is considered as a one-dimensional subspace of $H_{i}\subset H$. Let $p_{x}\in\mathcal{P}(H)$ be the projection with the range $\mathcal{E}_{x}$. The corresponding map $p\colon X\to\mathcal{P}(H)$ is strongly (but not uniformly) continuous, so the subspaces $\mathcal{E}_{x}\subset H$ give rise to the subbundle (in our sense) $\mathcal{E}$ of $H_{X}$. By [DD, Theorem 6], every section of $\mathcal{E}$ vanishes at some point. Therefore, $\mathcal{E}$ is not a locally trivial Hilbert bundle (since every locally trivial Hilbert bundle over a paracompact base space is trivial and thus have non-vanishing sections).
Moreover, $\mathcal{E}$ looks the same near each point of $X$ (more precisely, the group $\prod_{i=1}^{\infty}\mathcal{U}(H_{i})$ acts transitively on $X$ and this action lifts to the action on $\mathcal{H}$ by bundle automorphisms preserving $\mathcal{E}$), so a local triviality of $\mathcal{E}$ near one point would imply local triviality over the whole $X$. Therefore, for every open subset $X^{\prime}\subset X$, $\mathcal{E}$ is not trivial over $X^{\prime}$ and thus has no adapted trivializations over $X^{\prime}$.
Let $(c_{i})_{i\in\mathbb{N}}$ be a sequence of positive numbers converging to $\infty$. Then
$A=\oplus_{i=1}^{\infty}c_{i}\cdot\operatorname{Id}_{H_{i}}$
is a positive regular operator on $H$ with compact resolvents. The symmetry $r_{x}=2p_{x}-1$ commutes with $A$ and $\mathcal{A}_{x}=A\cdot r_{x}$ is an invertible self-adjoint operator with compact resolvents, which is neither essentially positive nor essentially negative. Since $r_{x}$ is a strongly continuous family of unitary operators, the family $\mathcal{A}_{x}$ is graph continuous by Proposition 2.2.
The positive spectral projection of $\mathcal{A}_{x}$ is $p_{x}$. By Proposition 7.2, each local trivialization adapted to $p$ is also adapted to $\mathcal{A}$. Since $p$ has no adapted trivializations, the family $\mathcal{A}$ has no adapted trivializations as well. Moreover, for every open subset $X^{\prime}\subset X$, the restriction of $\mathcal{A}$ to $X^{\prime}$ has no adapted trivialization.
@@
If one multiplies the family $\mathcal{A}_{x}$ from this example by the strongly continuous family of unitary operators $r_{x}$, then the resulting family $x\mapsto\mathcal{A}_{x}r_{x}=r_{x}\mathcal{A}_{x}=A$ will be Riesz continuous (and even constant). However, for every strongly continuous family of unitary operators $v_{x}\in\mathcal{U}(H)$, the family $x\mapsto v_{x}\mathcal{A}_{x}v_{x}^{*}$ is Riesz discontinuous at every point $x\in X$. This demonstrates the difference between the self-adjoint and non-self-adjoint cases, that is, between one-sided multiplication and conjugation by unitary operators, which we discussed in the Introduction.